40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Mar 5 06:27:14 2013, yuta, Update, LSC, intra-cavity power dependence on mirror misalignment PRM_PIT_20130305a.pngPRM_YAW_20130305b.pngITMX_YAW_20130305.pngITMY_YAW_20130305.png
    Reply  Tue Mar 5 23:00:08 2013, yuta, Update, LSC, Yarm and PRC g-factor from misalignment measurement ITMY_YAW_20130305_DC.pngPRM_PIT_20130305a_DC.pngPRM_YAW_20130305b_DC.png
       Reply  Tue Mar 5 23:37:11 2013, yuta, Update, SUS, PRM angular motion spectra PRMangularmotion.png
    Reply  Wed Mar 6 02:46:20 2013, Manasa, Update, LSC, PRMI locking for g-factor measurement POPDCspec_PRMIlocked.png
       Reply  Wed Mar 6 09:44:29 2013, Koji, Update, LSC, PRMI locking for g-factor measurement 
       Reply  Fri Mar 8 21:00:05 2013, Manasa, Update, PSL, PMC fixed 
Message ID: 8237     Entry time: Wed Mar 6 02:46:20 2013     In reply to: 8230     Reply to this: 8239   8263
Author: Manasa 
Type: Update 
Category: LSC 
Subject: PRMI locking for g-factor measurement 

 [Yuta, Manasa]

PRMI alignment procedure for carrier locking has been kept the same except that a couple of issues that have persisted are now taken care of.

We were able to keep PRMI locked for over a minute (POPDC measures 2200) .

1. Trigger levels to MICH and PRCL for PRMI locking have been changed
Whenever we enabled LSC controls to lock PRMI, ITMY moves haphazard if PRM is not aligned. This is because of the low trigger levels of POPDC which keeps MICH triggered all the time while we align PRM. Increasing POPDC trigger (Upper level : 1000 and Lower level:20) for both MICH and PRCL solved this problem and resulted in a more stable ITMY. Also this has stabilized locking greatly if the alignment is fair enough.

2. C1:SUS-ITMY_LSC_GAIN reset
Quoting Yuta's elog "  I found C1:SUS-ITMY_LSC_GAIN is somehow set to be 2.895 recently. I think this should be 1.0. Maybe this is why we had actuation imbalance in ITMs(elog #8212)."
This has been reset to 1.0 and it has not affected PRMI locking.


Mystery

1. Filter module (FM1) on PRCL and MICH show significant delay while enabling and disabling.

2. I tried to fix PMC alignment (PMC trans was 0.76). I was not able to get PMC trans more than 0.79.
PMC has been this way since yesterday.

3. MICH is still bright when locked (ASDC_OUT reads 0.08 for dark and 2.0 for bright). We suspect it is because of the AS55_I error offset that persists even after running LSCoffsets script.

4. PRMI shows some dither at 3Hz when locked.
POPDCspec_PRMIlocked.png

ELOG V3.1.3-